OpenStudy (anonymous):

Solve x2 + 4x + 8 = 0

OpenStudy (anonymous):

no solution imaginary my source= http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (anonymous):

\[x _{1,2}= -2\pm \sqrt{4-8}\] \[x_{1}=-2-2i\] \[x_{2}=-2+2i\] Hope this helps //Ali

OpenStudy (anonymous):

by defenition i^2=-1

jhonyy9 (jhonyy9):

x_1,_2=(-4+/- sqrt16-32))/2=(-4+/- 4i)/2= -2+/-2i or -2(1+/

jhonyy9 (jhonyy9):

= -2(1+/-i)